Sue is three times Mike's age. In four years, Sue will be two years older than twice Mike's age then. How old is each?
Answer by ikleyn(52788)

.
Sue is three times Mike's age. In four years, Sue will be two years older than twice Mike's age then. How old is each?
-----------------------------------------------------------

Your system of equations is

s     = 3m,           (1)
s + 4 = 2 + 2(m+4).   (2)

Substitute (1) into (2). You will get

3m + 4 = 2m + 8 + 2,

m = 8 + 2 - 4 = 6.

Answer. Mike is 6 years, Sue is 18 years now.